OpenStudy (victoriasushchik):

What is the next number in the Fibonacci sequence: 0, 1, 1, 2, 3, 5, 8, 13, 21, 34, ?

OpenStudy (anonymous):

Well, you can see that: 0+1=1 1+1=2 1+2=3 3+5=8 5+8=13 ... and so on, satisfying, \(a_{n-1}+a_{n}=a_{n+1}\)

OpenStudy (victoriasushchik):

i think it is 55 am i right??

OpenStudy (anonymous):

Yes 55 is right

OpenStudy (victoriasushchik):

thankyou very much

OpenStudy (anonymous):

no problem ...